OpenStudy (samsan9):

Describe the nature of the roots for this equation 2x^2+5x-7=0. A.) Two real, irrational roots B.) One real double root C.) Two real rational roots D.) Two complex roots

OpenStudy (anonymous):

First get the decremental \[\sqrt{b ^{2}-4ac}\] d is the inside of the root. if it was d = 0 then one real solution. d<0 2 imaginary d>0 2 real

OpenStudy (samsan9):

Sothe answer is D?

OpenStudy (anonymous):

as the rule is \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (samsan9):

it was C :(

OpenStudy (anonymous):

yea it is -7 not 7

OpenStudy (anonymous):

so you get d = 81

OpenStudy (samsan9):

i got -31...

OpenStudy (anonymous):

You substituted wrong :P  -*- = +